1) If -3 < k < 3, which of the following must be true?
2) If f is a linear function and f(3) = 6 and f(5) = 12, what is the slope of the graph of f?

Answer by stanbon(75887)

1) If -3 < k < 3, which of the following must be true?
---
Comment: data missing
------------------------------------
2) If f is a linear function and f(3) = 6 and f(5) = 12, what is the slope of the graph of f?
---
slope = (12-6)/(5-3) = 6/2 = 3
====================================
Cheers,
Stan H.
==========